House of Haddock has 5,000 shares outstanding and the stock price is $140. The company is expected to pay a dividend of $20 per share next year and thereafter the dividend is expected to grow indefinitely by 5% a year. The President, George Mullet, now makes a surprise announcement: He says that the company will henceforth distribute half the cash in the form of dividends and the remainder will be used to repurchase stock. The repurchased stock will not be entitled to the dividend.

a-1. What is the total value of the company before the announcement?

Total value            $

a-2. What is the total value of the company after the announcement?

Total value            $   

a-3. What is the value of one share?

Value per share            $

b. What is the expected stream of dividends per share for an investor who plans to retain his shares rather than sell them back to the company? Check your estimate of share value by discounting this stream of dividends per share. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Year Expected Dividend
1 $
2 $
Answer #1

a-1). Value of company before announcement = 5000 * 140

= $ 700,000.

a-2). Value of company after announcement = 700,000 + 5 % of 700,000

= 700000 + 35000

= $ 735000.

a-3). Value per share = $ 20. (Value per share will be equal to dividend per share only).